OpenStudy (anonymous):

write the equation y-4=-2(x-5) in slope-intercept form y=-2x+14 y=-2x+5 y=-2x+9 y=-2x+6

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

To find what you're looking for. y-4=-2(x-5) First distribute  You get. y-4=-2x+10 Now add for on both sides. y=-2x+14